# Collaborating with AI Agents: A Field Experiment on Teamwork, Productivity, and Performance

### v2

@@ 2 Methodology
Our study was preregistered and approved by the MIT Committee on the Use of Humans as Experimental Subjects. [^1] See osf.io/jfzha and osf.io/95dhu. Any non-pre-registered analyses are labeled post hoc.
@@ 2.1 The Pairit platform
Once a participant enters the Pairit (Figure 1) platform, they are randomized into a queue for either the human-human or the human-AI condition. In the human-human condition, the participant collaborates with another human participant. In the human-AI condition, the participant collaborates with an AI agent. In addition to team-level randomization, the AI models are also configured to accept prompt and fine-tuning randomization. Pairit’s left panel is the task panel in which participants can participate in the task work process (in the context of this experiment to create, edit, and submit ads). The platform includes a carousel of images participants can choose from, and the participants (as well as the AI) can also access the Dall-E 3 application programming interface (API) to generate new images. The ad copy creation and editing panel includes a headline, primary text, and a description. All of the edits—including the image selection and image generation, as well as the ad copy edits—are synchronized in real-time across the participants, whether they are human or AI. On the right is the chat panel in which participants can chat in real time with either another human participant or an AI agent. The participants can submit the ads and the interface will reset so that they can work on and then submit the next ad. To the best of our knowledge, Pairit is the first platform to faciliate real-time collaboration between human-human or human-AI teams with team- and model-level randomization and real-time chat and editing on synchronized text and image interfaces in the context of multimodal human-AI collaboration workspaces.
Figure 1: The Pairit platform. On the left is the task panel, and on the right is the chat panel. In the Human-Human condition, chat messages and edits on the task panel, including text edits, image selections, and AI image generations, are synchronized in real-time. In the Human-AI condition, the participant chats with an AI agent with full context of the user interface (UI; see Section 2.2), and the AI can edit text, select images, and generate AI images.
@@ 2.2 AI agent
The platform uses OpenAI’s multimodal gpt-4o model (specifically gpt-4o-2024-08-06), which processes text and images in a single model. To give the AI full context of the task and collaboration, each API call to gpt-4o is prompted with the information on the screen so it has the full context of the collaboration. In the prompt, we include the same text of the task given to participants, previous submissions, current copy, elapsed time, history of its own actions, history of its chain-of-thought, chat history, and general instructions. Moreover, a screenshot is taken of the image after every change and included as an input so the AI can observe the image and its evolution over time and edits. These details enable human collaboration with a fully functional, multimodal AI agent.
@@
To ensure the agent is taking actions appropriately, we use chain-of-thought prompting (Wei et al., 2023). [^2] We used OpenAI’s structured output feature for chain-of-thought prompting. See their documentation. Specifically, we prompt the model with questions to reflect on the state of the collaboration. These questions were necessary to avoid undesirable model behavior, such as repeatedly sending the same message and were developed and perfected through trial and error.
@@ 2.3 Procedure
After the completion of the ad creation task, the participants answered a 35-item teamwork quality survey (Hoegl and Gemuenden, 2001). The original survey consisted of 38 items, each on a 7-point Likert scale; however, we removed 3 items from the communication facet that did not apply to a single-session online collaboration task. In addition, we asked four questions regarding participants’ perception of AI, all on a 7-point Likert scale. These included two questions about their experience using AI (i.e., “I have used artificial intelligence (AI) chatbots before (e.g., ChatGPT, Bard).” and “I had a positive experience using AI chatbots.”), one question about their perception of their partner as an AI (i.e., “I believe my partner was an AI during the task.”) and a final question in which we revealed the identity of their partner (human or AI) and ask whether that changed the perception of the quality of the collaboration (i.e., “Your partner was [an AI assistant/a human]. Knowing this, to what extent has your perception of the quality of your collaboration changed?”). 85% of participants completed the post-task survey.
@@ 2.4 Participants
We pre-registered a sample of 2,500 participants from Prolific (www.prolific.com) based in the US, with representative stratification across gender and ethnicity. In total, 2310 participants completed the task. Of the 2500 participants who entered, 23 did not enter the matching queue (see Section 2.3 for details) and were removed from the study. A further 167 participants quit the study or were timed out before they were matched to a partner. The overall attrition rate was 7.6%. The study was conducted from October 15 to 18, 2024, and took a median of 46.32 minutes to complete. The participants were paid \$9.
@@ 2.5 Summary statistics and randomization checks
In total, the dataset includes 2,310 participants, 1,834 teams, 11,138 display ad submissions, 183,691 messages, 63,656 edits on images, 1,960,095 edits on ad copy, and 10,375 AI-generated images. To ensure that the randomization procedure successfully balanced covariates across experimental conditions, Table 1 compares key participant characteristics between the human-human and human-AI conditions. These include demographic variables (e.g. age, gender), participants’ work status, and the Big Five personality traits. No significant differences were detected apart from a very small difference in part-time employment status across treatment conditions, indicating broad-based, balanced randomization across participants.
@@
Individuals 2310 1258 1052 - -
Teams 1834 1258 576 - -
Gender (% Male) 50.8% 50.4% 51.3% -0.401 (0.021) 0.688
Age 42 $\pm$ 14 43 $\pm$ 15 42 $\pm$ 14 0.987 (0.599) 0.324
Full-Time 47.8% 48.5% 47.1% 0.688 (0.021) 0.491
Part-Time 14.7% 13.4% 16.3% -2.009 (0.015) 0.045
Data expired 13.4% 13.2% 13.6% -0.279 (0.014) 0.780
Not in paid work 10.3% 10.9% 9.7% 0.943 (0.013) 0.346
Unemployed (and job seeking) 8.8% 9.1% 8.6% 0.428 (0.012) 0.668
Other 4.2% 4.2% 4.3% -0.077 (0.008) 0.939
Start new job in a month 0.6% 0.8% 0.5% 0.974 (0.003) 0.330
Openness 0.71 $\pm$ 0.20 0.71 $\pm$ 0.21 0.72 $\pm$ 0.19 -0.854 (0.008) 0.393
Conscientiousness 0.79 $\pm$ 0.18 0.78 $\pm$ 0.18 0.79 $\pm$ 0.17 -1.166 (0.007) 0.244
Extraversion 0.56 $\pm$ 0.21 0.55 $\pm$ 0.21 0.56 $\pm$ 0.21 -0.962 (0.009) 0.336
Agreeableness 0.70 $\pm$ 0.22 0.70 $\pm$ 0.21 0.69 $\pm$ 0.24 0.534 (0.009) 0.594
Neuroticism 0.48 $\pm$ 0.23 0.49 $\pm$ 0.24 0.48 $\pm$ 0.22 1.538 (0.010) 0.124
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Personality traits are normalized from a 7-point Likert scale. For t-statistics, gender is coded as 1 for male and 0 for female, and employment status is 1 for the listed category (e.g., full-time) and 0 otherwise.
@@ 2.6 Incentives
To incentivize participants to create high-quality ads, we informed them that they were eligible for additional \$100 prizes based on the quantity and quality of the ads they submitted, as well as the performance of the ads in the marketing field experiment. Participants were explicitly instructed that “the greater the number of ads, the greater your chances—but not if the ads are of low quality." Ultimately, two participants were awarded \$100 each for producing the greatest number of high-performing ads.
@@ 2.7 Message labeling
To label messages into categories and intent, we prompt gpt-4o-mini-2024-07-18 for each message independently and ask for a label for a category and an intent of the message. We enforce that the labels are from the set of pre-determined labels using OpenAI’s Structured Outputs API.
2.8 AI evaluation of ad quality
Ad mockups.
To obtain ratings of display ads that are as close as possible to display ads one would see on digital ad publishers, we first created mockups of each display ad. An ad mockup is a simulated representation of a display ad designed to closely resemble the appearance and functionality of an actual ad as it would appear on digital ad publishers’ platforms. We built a web application that populates the image, ad copy, shortened link, call-to-action, and other user interface items—including the Like, Comment, Share, and Close buttons, the profile picture, and the Sponsored tag. Screenshots were then programmatically taken of each mockup. See Figure 3 for examples.
Figure 3: Examples of ad mockups.
AI ratings.
If AI ratings are shown to predict human evaluations of ad quality and field evaluations of ad performance, they can potentially provide cost-effective alternatives to ad testing. To obtain AI ratings of ad quality, we prompt OpenAI’s multimodal gpt-4o-mini-2024-07-18, which supports image input and structured outputs. [^3] See OpenAI’s documentation on vision and structured outputs. To make AI evaluations comparable to human evaluations, we asked the same questions of the AI as those given to human evaluators (see Section 2.9). Each item was evaluated on a 7-point Likert scale. The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad."
2.9 Human evaluation of ad quality
@@
We used the same mockups of display ads as used for AI evaluation of ad quality, as explained in Section 2.8. For each ad, we asked participants to provide three ratings regarding the quality of the text, the image, and the estimated click through rate. Each item was rated on a 7-point Likert scale. The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad." See Figure 4 for an example of an item on the survey.
Figure 4: The user interface for the ad quality survey.
@@ 2.10 Field evaluation of ad performance
To run as many ads as possible while sampling evenly from human-human and human-AI teams and from the human predictions of ad quality, we implemented a stratified sampling approach. We obtained a sample of 2,000 ads from a total of 11,138 ads, sampling between one and two ads from each of the 1,834 teams. We first divided the ads into those created by human-human teams and human-AI teams and then further divided the ads into 10 strata according to the human predictions of the likelihood of user clicks, from low to high predicted click probability (see Section 2.9). We removed 8 ads that potentially violated moderation policies (due to the inclusion of e.g. violence, sexual, or drug content). We controlled for advertising spend in our analyses to account for auto-bidding.
To prevent overlapping audience targeting, we assigned five unique ads to each of 400 campaigns, structuring them as 5-ad split tests within individual campaigns (noting the platform’s limit of five split tests per campaign). To further ensure no overlap across campaigns, we allocated a random set of 133 ZIP codes to each campaign, selecting ZIP codes with populations between 10,000 and 100,000.555The data is from https://data.census.gov/table/ACSDT5Y2020.B01003. We tested the robustness of this allocation across ZIP codes by conducting one-way ANOVA on population and income, yielding non-significant results: $F(399,53199)=0.954$, $p=0.734$ for population, and $F(399,53199)=0.973$, $p=0.636$ for income.
@@ 2.11 Model specifications
where $Y_{t}$ represents the outcome for team $t$, HAIi is 1 if individual $i$ is working with an AI and 0 if individual $i$ is working with another participant, $X_{i}$ includes demographic controls including age, gender, and Big Five personality traits, and $\epsilon_{i}$ is the error term. The outcomes of interest $Y_{t}$ include the number of interactions on the platform (i.e., messages, copy edits, image selects, AI image generations, submissions), fraction of messages by category, quality scores, and fraction of copy completed. For the analysis of ad quality, we perform analysis on the rating level where $i$ is an ad rating. For team-level analysis, we use the following model:
$Y_{t}=\delta_{t}\text{HAI}_{t}+\beta X_{t}+\epsilon_{t}\text{,}$ (2)
where the outcome $Y_{t}$ is the same as above, HAIt (Human-AI) is 1 if the team was Human-AI and 0 if the team was Human-Human, $X_{t}$ includes team-level demographic controls including average age, gender composition, and average Big Five personality traits, and $\epsilon_{t}$ is the error term.
@@
$Y_{tc}=$ $\ \delta_{t}\text{HAI}_{tc}+\text{Text}_{tc}+\text{Image}_{tc}+\text{Click}_{tc}$
$+\text{Spend}_{tc}+u_{c}+\epsilon_{tc}\text{,}$ (3)
@@ 3 Results
3.1 Work Processes
Collaborating with AI increases communication and decreases direct engagement with work product.
To test how human-AI collaboration affects communication, we measured the number of messages sent by participants in both human-human and human-AI groups. Participants worked collaboratively with and sent messages to their partners in real-time, for both human-human and human-AI groups. In the human-human group, participants messaged their human partners; in the human-AI group, participants messaged the AI. The Pairit platform logs a timestamped record of each message sent by a participant or AI.
Participants communicated significantly more in human-AI teams than in human-human teams. Participants in human-AI teams sent 63% more messages than those in human-human teams, with consistent results when demographic covariates are included (Table 2). These results indicate that collaborating with an AI partner encourages more frequent communication.
Messages
Intercept 21.504∗∗∗ 9.977∗∗
(0.484) (2.999)
Human-AI 13.669∗∗∗ 13.671∗∗∗
(0.834) (0.834)
Demographics No Yes
Observations 2310 2310
Table 2: Collaborating with AI increases communication at the individual level.
On the other hand, participants in Human-AI teams made 71% fewer direct copy edits compared to Human-Human teams (Table 3). The results on copy edits and messages suggest that collaboration with AI shifted participants’ focus from generating text to instructing and suggesting the AI to generate text, as we will see in greater detail in the next set of analyses. A post hoc analysis of the correlation between user actions found that the number of messages sent and copy edits were negatively correlated ($R^{2}=-0.17$, $p=3.5^{-17}$), further supporting this interpretation. However, participants in the Human-AI condition still engaged in direct editing, indicating that while AI collaboration reduced the frequency of manual edits, it did not eliminate them entirely. Unlike previous studies, such as Chen and Chan (2024), which artificially constrained participants to specific modalities when interacting with AI, our results show that in more realistic collaborative settings, where humans and agents can freely allocate effort to different tasks, participants utilized both direct engagement with work product (ad editing) and interaction through messaging.
Copy Edits Image Selects AI-Generated Images
Intercept 1367.225∗∗∗ 1630.869∗∗∗ 23.219∗∗∗ 23.447∗∗∗ 4.234∗∗∗ 5.921∗∗∗
(30.059) (117.896) (0.579) (3.727) (0.149) (0.955)
Human-AI -974.224∗∗∗ -971.857∗∗∗ 7.432∗∗∗ 7.403∗∗∗ 0.401 0.420
(32.939) (32.812) (0.907) (0.911) (0.214) (0.212)
Demographics No Yes No Yes No Yes
Observations 2310 2310 2310 2310 2310 2310
Table 3: Collaborating with AI reduces copy edits but increases image edits and AI image generations at the individual level.
Human-AI teams focus on content and process over social and emotional communication.
In addition to the number of messages sent across conditions, we investigated whether the content of messages varied across conditions. We used gpt-4o-mini to label each message independently with one of five message categories: Process, Content, Social, Emotional, Feedback, and Other (see Section 2.7 for more details). In our analysis, we found that both the AI and the humans in human-AI teams sent 18% (i.e., (0.4060.345)/0.345=0.177(0.406-0.345)/0.345=0.177) more content- and process-oriented messages, while human-human teams sent 29% (i.e., (0.5790.449)/0.449=0.29(0.579-0.449)/0.449=0.29) more social and emotional messages (Figure 5, Table 4). This shift indicates that collaboration with AI emphasizes task-related communication over social interaction, possibly because participants can focus more on the task without needing to navigate the social or emotional aspects of collaboration.
Figure 5: Participants in human-AI teams send more process- and content-related messages while those in human-human teams send more social and emotional messages.
To further examine how communication differs across conditions, we labeled the messages with one of 36 message intentions. These intentions capture a range of functions, such as suggestions, instructions, and prioritization, as well as rapport building and self-assessment. We found notable differences in the distribution of message intentions between the human-human and human-AI groups in each message category (see Table 4 and Figure 6). For example, in the Content category, human-human teams send more messages that exhibit confusion and clarification, and human-AI teams send more messages about brainstorming, confirmation, acknowledgment, suggestion, agreement, instruction, and judgment. The results are similar for process-related messages. For messages in the social and emotional categories, human-human teams send more apologies and messages about humor and concern, while human-AI teams send more suggestions and messages about appreciation, motivation, confirmation, and satisfaction. These findings further support the idea that human-AI collaboration prioritizes task-oriented communication, while human-human teams engage more in social and emotional exchanges.
Content Process Social Emotional Feedback Other
Intercept 0.156∗∗∗ 0.189∗∗∗ 0.415∗∗∗ 0.164∗∗∗ 0.053∗∗∗ 0.051∗∗∗
(0.025) (0.021) (0.028) (0.017) (0.010) (0.011)
Human-AI 0.036∗∗∗ 0.025∗∗∗ -0.085∗∗∗ -0.045∗∗∗ 0.019∗∗∗ 0.018∗∗∗
(0.006) (0.006) (0.007) (0.005) (0.003) (0.003)
Demographics Yes Yes Yes Yes Yes Yes
Observations 2310 2310 2310 2310 2310 2310
  • Notes: p<<0.05; ∗∗p<<0.01; ∗∗∗p<<0.001. Heteroskedasticity-robust standard errors are reported.
Table 4: Collaborating with AI increases content-, process-, and feedback-related messages while decreasing social and emotional messages.
Figure 6: Message labeled by intent for each category.
Survey reports showed no differences in perceptions of teamwork quality.
Despite the substantial differences observed in communication patterns and task-focused behavior between human-human and human-AI teams, survey responses indicated no significant differences in participants’ perceptions of teamwork quality across conditions. This lack of differentiation might stem from the way participants perceive AI: unlike human collaborators, AI agents may not evoke the same social or emotional expectations. These results suggest that while collaboration dynamics differ, participants may not attribute those differences to changes in teamwork quality when working with AI, though teammate identity has been found not to affect trust (Zhang et al., 2023). For future studies relying on survey responses, this highlights a potential limitation: actual differences in collaboration dynamics may not be fully reflected in self-reported measures of teamwork quality.
3.2 Productivity
Having established that collaborating with AI agents allows participants to focus on the task without the social coordination costs typically associated with human collaboration, we now examine its impact on productivity. Specifically, we analyze the number of submissions per team and individual and the completion rate of ad copy.
Human-AI teams produce a similar amount of output with half as many workers.

### v3

@@ 3 Methodology
Our study was preregistered and approved by the MIT Committee on the Use of Humans as Experimental Subjects. [^1] See osf.io/jfzha and osf.io/95dhu.
@@ 3.1 The Pairit platform
Once a participant enters the Pairit (Figure 1) platform, they are randomized into a queue for either the human-human or the human-AI condition. In the human-human condition, the participant collaborates with another human participant. In the human-AI condition, the participant collaborates with an AI agent. Pairit’s left panel is the task panel in which participants can participate in the task work process (in the context of this experiment to create, edit, and submit ads). The platform includes a carousel of seven pre-selected stock images provided by the think tank that the participants and the AI agent can choose from, and the participants (as well as the AI) can also access the Dall-E 3 application programming interface (API) to generate new images. The ad copy creation and editing panel includes a headline, primary text, and a description. All edits, including image selection, image generation, and ad copy changes, are synchronized in real-time across participants. The chat panel displays a “Typing…” indicator when the partner is composing a message. When participants submit an ad, the canvas clears and resets for the next submission. On the right is the chat panel in which participants can chat in real time with either another human participant or an AI agent. To the best of our knowledge, Pairit is the first platform to facilitate real-time collaboration between human-human or human-AI teams with team- and model-level randomization and real-time chat and editing on synchronized text and image interfaces in the context of multimodal human-AI collaboration workspaces.
Figure 1: The Pairit platform. On the left is the task panel, and on the right is the chat panel. In the human-human condition, chat messages and edits on the task panel (including text edits, image selections, and AI image generations) are synchronized in real-time. In the human-AI condition, the participant chats with an AI agent with full context of the user interface (UI; see Section 3.2), and the AI can edit text, select images, and generate AI images.
@@ 3.2 AI agent
The platform uses OpenAI’s multimodal gpt-4o model (specifically gpt-4o-2024-08-06), which processes text and images in a single model. To give the AI full context of the task and collaboration, each API call to gpt-4o is prompted with the information on the screen so it has the full context of the collaboration. In the prompt, we include the same text of the task given to participants, previous submissions, current copy, elapsed time, history of its own actions, history of its chain-of-thought, chat history, and general instructions. Moreover, a screenshot is taken of the image after every change and included as an input so the AI can observe the image and its evolution over time and edits. These details enable human collaboration with a fully functional, multimodal AI agent. The prompt template is shown in Appendix K.
@@
To ensure the agent is taking actions appropriately, we used chain-of-thought prompting (Wei et al., 2023). [^2] We used OpenAI’s structured output feature for chain-of-thought prompting. Specifically, we prompt the model with questions to reflect on the state of the collaboration. These questions were necessary to avoid undesirable model behavior, such as repeatedly sending the same message and were developed and perfected through trial and error.
@@ 3.3 Procedure
After the completion of the ad creation task, the participants answered questions regarding their perception of AI, all on a 7-point Likert scale. These included two questions about their experience using AI (i.e., “I have used artificial intelligence (AI) chatbots before (e.g., ChatGPT, Bard).” and “I had a positive experience using AI chatbots.”), one question about their perception of their partner as an AI (i.e., “I believe my partner was an AI during the task.”) and a final question in which we revealed the identity of their partner (human or AI) and ask whether that changed the perception of the quality of the collaboration (i.e., “Your partner was [an AI assistant/a human]. Knowing this, to what extent has your perception of the quality of your collaboration changed?”). 85% of participants completed the post-task survey.
@@ 3.4 Participants
We pre-registered a sample of 2,500 participants from Prolific (www.prolific.com) based in the US, with representative stratification across gender and ethnicity. In total, 2,310 participants completed the task. Of the 2,500 participants who entered, 23 did not enter the matching queue (see Section 3.3 for details) and were removed from the study. A further 167 participants quit the study or were timed out before they were matched to a partner. In 83 human-human sessions (14% of HH teams), the matched partner did not complete the task due to dropout or timeout. Because these sessions do not represent collaborative work and submissions required confirmation by both participants, we exclude them from all analyses, yielding a final sample of 2,234 participants (1,258 human-AI, 976 human-human). The overall attrition rate was 8.9%. The study was conducted from October 15 to 18, 2024, and took a median of 46.32 minutes to complete. The participants were paid \$9.
@@ 3.5 Summary statistics and randomization checks
In total, the dataset includes 2,234 participants, 1,751 teams, 11,024 display ad submissions, 182,607 messages, 62,119 edits on images, 1,889,559 edits on ad copy, and 10,074 AI-generated images. To ensure that the randomization procedure successfully balanced covariates across experimental conditions, Table 1 compares key participant characteristics between the human-human and human-AI conditions. These include demographic variables (e.g. age, gender), participants’ work status, and the Big Five personality traits. No significant differences were detected across treatment conditions, indicating balanced randomization across participants.
@@
Individuals 2234 1258 976
Teams 1751 1258 493
Gender (% Male) 50.8% 50.4% 51.3% -0.438 (0.021) 0.661
Age 42 $\pm$ 14 43 $\pm$ 15 42 $\pm$ 14 1.296 (0.608) 0.195
Full-Time 47.7% 48.5% 46.7% 0.830 (0.021) 0.407
Part-Time 14.6% 13.4% 16.2% -1.883 (0.015) 0.060
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. For t-statistics, gender is coded as 1 for male and 0 for female, and employment status is 1 for the listed category. Balance checks for Big Five personality traits appear in Table A.
@@ 3.6 Incentives
To incentivize participants to create high-quality ads, we informed them that they were eligible for additional \$100 prizes based on the quantity and quality of the ads they submitted, as well as the performance of the ads in the marketing field experiment. This incentive structure was identical across human-human and human-AI conditions. Participants were explicitly instructed that “the greater the number of ads, the greater your chances, but not if the ads are of low quality.” Ultimately, two participants were awarded \$100 each.
@@ 3.7 Message labeling
To label messages into categories, we prompt gpt-4o-mini-2024-07-18 for each message independently and ask for a category label. We enforce that the labels are from the set of pre-determined labels using OpenAI’s Structured Outputs API. Categories include content, process, social, emotional, feedback, and other. We group content and process messages as task-oriented communication, and social and emotional messages as interpersonal communication. See Appendix J for the prompts used in this analysis.
3.8 Evaluation of ad quality
@@
Ad mockups.
To obtain ratings of display ads that are as close as possible to display ads one would see on digital ad publishers, we first created mockups of each display ad. An ad mockup is a simulated representation of a display ad designed to closely resemble the appearance and functionality of an actual ad as it would appear on digital ad publishers’ platforms. We built a web application that populates the image, ad copy, shortened link, call-to-action, and other user interface items, including the Like, Comment, Share, and Close buttons, the profile picture, and the Sponsored tag. Screenshots were then programmatically taken of each mockup.
@@
We used mockups of display ads. For each ad, we asked participants to provide three ratings regarding the quality of the text, the image, and the estimated click through rate. Each item was rated on a 7-point Likert scale. The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad." See Figure F for an example of an item on the survey.
@@ 3.9 Field evaluation of ad performance
To run as many ads as possible while sampling evenly from human-human and human-AI teams and from the human predictions of ad quality, we implemented a stratified sampling approach. We obtained a sample of 2,000 ads from a total of 11,138 ads, sampling between one and two ads from each of the 1,751 teams. We first divided the ads into those created by human-human teams and human-AI teams and then further divided the ads into 10 strata according to the human predictions of the likelihood of user clicks, from low to high predicted click probability (see Section 3.8). We removed 8 ads that potentially violated moderation policies (due to the inclusion of e.g. violence, sexual, or drug content). We controlled for advertising spend in our analyses to account for auto-bidding. Consistent with other analyses, field analyses exclude 37 ads from human-human sessions where one partner did not complete the collaboration.
To prevent overlapping audience targeting, we assigned five unique ads to each of 400 campaigns, structuring them as 5-ad split tests within individual campaigns (noting the platform’s limit of five split tests per campaign). To further ensure no overlap across campaigns, we allocated a random set of 133 ZIP codes to each campaign, selecting ZIP codes with populations between 10,000 and 100,000.444The data is from https://data.census.gov/table/ACSDT5Y2020.B01003. We tested the robustness of this allocation across ZIP codes by conducting one-way ANOVA on population and income, yielding non-significant results: $F(399,53199)=0.954$, $p=0.734$ for population, and $F(399,53199)=0.973$, $p=0.636$ for income.
@@ 3.10 Model specifications
where $Y_{i}$ is the outcome, HAIi is 1 if the unit is in the human-AI condition and 0 otherwise, $X_{i}$ includes demographic controls (age, gender, Big Five personality traits), and $\epsilon_{i}$ is the error term. The unit of analysis $i$ varies by outcome: individual-level for messages, copy edits, image edits, AI image generations, submissions, message categories, and delegation rates; team-level for team submissions; and rating-level for ad quality scores. For team-level analyses, $X_{i}$ includes team-level controls (average age, gender composition, average personality traits).
@@
$Y_{tc}=$ $\ \delta_{t}\text{HAI}_{tc}+\text{Text}_{tc}+\text{Image}_{tc}+\text{Click}_{tc}$ (2)
$+\text{Spend}_{tc}+u_{c}+\epsilon_{tc}\text{,}$
@@
Mediation analysis.
To trace the pathway from treatment to communication to production actions to quality outcomes, we estimate sequential regressions using Equation 1, substituting the relevant predictor and outcome at each step. For example, we regress production actions (i.e., delegation, submissions, manual edits) on communication patterns (i.e., task-oriented, interpersonal), and quality outcomes on delegation. This approach identifies which communication patterns are associated with production behaviors, and how delegation relates to quality outcomes.
Recognition of partner identity.
To test whether correct identification of partner type moderates treatment effects, we estimate:
$Y_{i}=$ $\beta_{0}+\beta_{1}\text{HAI}_{i}+\beta_{2}\text{Recognition}_{i}$ (3)
$\quad+\beta_{3}(\text{HAI}_{i}\times\text{Recognition}_{i})+\gamma X_{i}+\varepsilon_{i}\text{.}$
A significant $\beta_{3}$ indicates that the treatment effect differs depending on whether participants correctly identified their partner type.
To construct a measure of recognition, we use the partner perception question (“I believe my partner was an AI during the task”), measured on a 7-point Likert scale from strongly disagree to strongly agree. Using a median split (score $\geq$ 4 = “believed AI”), we define Recognition = 1 if participants correctly identified their partner type: believed AI in the H-AI treatment, or believed human in the H-H treatment.
Delegation.
We measure delegation as the fraction of text editing work performed by the partner rather than the focal participant. We defined delegation as one minus the participant’s character contribution divided by the session’s total character contribution, which, on average for human-human teams, is mechanically 50%. For human-AI teams, we detect AI-generated content by identifying large text jumps (simultaneous edits of more than 10 characters), which is how the AI produced the content compared to the human editing incrementally.
Output diversity.
To measure output diversity, we embed each submission’s ad copy (headline, primary text, and description) using OpenAI’s text-embedding-3-small model. We compute the centroid (mean embedding vector) for each treatment group and calculate the cosine distance from each submission to its own treatment’s centroid. Higher values indicate greater distance from the typical output in that condition, reflecting more diverse or atypical content. We aggregate to the user level by averaging across submissions and estimate treatment effects using Equation 1. This analysis is restricted to users who submitted at least one ad with text content (NN=2,034 of 2,234 participants).
@@ 4 Results
4.1 Productivity, Quality and Output Diversity
We begin by examining the primary outcomes of human-AI collaboration: productivity, ad quality, and output diversity. These outcomes reveal a consistent pattern: human-AI teams achieve higher individual productivity and better text quality, but at the cost of reduced image quality and less diverse outputs.
Collaborating with AI agents increases individual-level productivity.